# Vertex reconstruction in the TAO experiment

### v1

@@ 1 Introduction
In the case of small-scale detectors like TAO, a compact detector design results in insufficient photon time-of-flight resolution and hinders the construction of precise time templates. Instead, CCA relies solely on the spatial distribution of charge to directly and efficiently compute the charge-weighted centroid with low algorithmic complexity and high computational efficiency, making it ideal for reliable and rapid reconstruction of large event sets. DLA eliminates the need for constructing templates by directly learning nonlinear mappings from the charge and time data, achieving exceptional reconstruction accuracy. Its inherent parallel processing capabilities enable large-scale data handling, making it a powerful complement to traditional approaches. Therefore, this paper focuses on CCA and DLA and systematically investigates the optimization strategies and performance characteristics of the two algorithms, proposing the corresponding high-precision vertex reconstruction solutions for the TAO experiment and ensuring robust preparation for the upcoming data acquisition phase.
@@ 2 The TAO detector
The CD configuration of the TAO experiment is depicted in Fig. 1. The innermost component of the CD is the liquid scintillator system, which consists of a spherical acrylic vessel (AV) with an inner diameter of 1.8 $\mathrm{m}$, filled with 2.8 tonnes of GdLS. To mitigate external radioactive and incomplete energy deposition near the GdLS boundary, a fiducial volume of 0.65 $\mathrm{m}$ radius is designated as the event selection criterion. A total of 4024 Silicon Photomultiplier (SiPM) tiles, each measuring $50.7\times 50.7~\mathrm{mm}^{2}$, are installed in an array configuration on the acrylic vessel surface, which is supported by a copper shell structure. Each tile consists of 32 SiPMs with active areas of $12\times 6~\mathrm{mm}^{2}$, and includes two readout channels, resulting in a total of 8048 readout channels. The SiPM tile coverage in the TAO central detector surpasses 93.5%. This high coverage substantially improves photon collection efficiency, thereby ensuring an energy resolution $\sim$ 2% at 1 MeV.
@@ 3.2 Optimizations
First, we consider a simple case that utilizes Eq. (1) to calculate the $R$ reconstruction results for electrons directly without electronic effects. Fig. 3 presents the reconstruction results for electrons with energies ranging from 1 MeV to 9 MeV, uniformly generated within the GdLS. It is observed that the reconstruction bias $B_{\text{reco}}$ becomes more pronounced near the dual-opening location, where $B_{\text{reco}}=R_{\text{reco}}-R_{\text{true}}$, with $R_{\text{reco}}$ and $R_{\text{true}}$ representing the reconstructed radius and the true radius of the energy deposition position, respectively. This is primarily due to an increase in photon leakage close to the dual-opening, resulting in a lower $R_{\text{reco}}$. To address the impact of the dual-opening effect, we propose the following dual-opening correction scheme:
@@
Fig. 6: Fitting curves of the two-dimensional distributions of $\mathit{R_{\text{true}}}$ and $\mathit{R_{\text{reco}}}$. The 137Cs curve (E = 0.662 MeV) is based on simulations along the calibration path, while the gamma ($\gamma$), positron ($e^{+}$), and electron ($e^{-}$) curves are derived from uniform simulations in the GdLS. Panels (a) and (b) show the results without and with DN correction, respectively.
@@ 4.2 The VGG-T model
The Visual Geometry Group (VGG) at the University of Oxford proposed a classic CNN model known as the VGG network architecture in 2014 Simonyan:2014cmh . This model is characterized by its deep architecture, which allows for extracting hierarchical features from input data. Its advantages include improved performance on various image classification tasks, a straightforward design that facilitates transfer learning, and robustness to different image resolutions. These features make the VGG network popular for many computer vision applications. Building upon the VGG framework, we designed the VGG-T model tailored to predict the vertex of TAO events.
@@
The predicted vertex coordinates are represented in the Cartesian space as $\mathbf{r}(x,y,z)$, with values constrained to the range [-900 mm, 900 mm]. To enforce this constraint, the loss function incorporates two penalty terms: $\text{Penalty}_{\text{upper}}$ for values exceeding 900 mm and $\text{Penalty}_{\text{lower}}$ for values below -900 mm. The total loss function is defined as follows:
@@ 4.3 The ResNet-T model
We utilize this technique to develop a deeper network architecture called ResNet-T, which is tailored for reconstruction tasks in the TAO experiment. As shown in Fig. 12, ResNet-T incorporates thirty-four convolutional layers beyond the original VGG architecture, resulting in a $54\%$ increase in the total number of parameters. This substantial increase in model capacity allows ResNet-T to learn more complex features and relationships within the data.
Throughout the training process, ResNet-T utilizes the same training settings as those applied to VGG-T, ensuring consistency while allowing us to explore the benefits of depth via the residual connections. This approach facilitates better performance in reconstruction tasks and fosters a more stable training process, ultimately contributing to improved accuracy and efficiency in the TAO experiment.
@@ 5.1 Performance improvement of optimized CCA
In Section 3.2, we introduced optimizations to CCA targeting dual-opening, $R$ reconstruction nonlinearity, and electronic effects in the TAO CD to enhance reconstruction performance. Since the improvements to the reconstruction of the angles $\theta$ and $\phi$ are negligible, this section is dedicated solely to systematically evaluating and discussing the impact of these optimizations on the $R$ reconstruction using two metrics: bias (mean of $R_{\rm rec}-R_{\rm true}$) and resolution (its standard deviation).
@@ 5.2 Comparison of reconstruction performance between CCA and DLA
The reconstruction results of DLA in $R$ reconstruction are shown in Fig. 16. In terms of bias, the outcomes for all three particle types are generally controlled within 1 mm. However, positrons with initial kinetic energies below 2 MeV exhibit a larger bias exceeding 1 mm. This phenomenon arises because the gamma rays produced during the annihilation of low-energy positrons carry a higher proportion of energy, leading to a more dispersed distribution of energy deposition positions and consequently increasing the reconstruction error. Regarding reconstruction resolution, the $R$ resolution values demonstrate a decreasing trend as the deposited energy increases. Among the particles, electrons exhibit the best resolution performance, benefiting from their more concentrated energy deposition under the same deposited energy conditions, which enables more precise vertex localization and reduces statistical errors. Comparing the results of the two network architectures, ResNet-T, with its deeper network layers, shows improvements in $R$ reconstruction bias and resolution, particularly notable resolution enhancement. It is worth noting that the improvement in electron resolution is the most significant, reaching up to 13%, highlighting the substantial optimization effect of this network structure on electron reconstruction performance.
@@
Fig. 17 compares electron reconstruction performance between the CCA and DLA, revealing several key advantages of the latter. The DLA yields smaller biases with more stable trends than the CCA and delivers markedly better resolution under all conditions. Although the DLA’s resolution slightly degrades within the fiducial volume (R<650R<650\,$m$), this reflects its superior handling of boundary events. These edge events, despite involving more complex electronic effects, benefit from a more concentrated charge distribution on the SiPM channels (as shown in Fig. 9(a)) and enhanced distinctiveness of image features, enabling the DLA to handle boundary events more effectively. This variation in performance suggests that DLA is particularly effective at processing spatially localized signal patterns that are characteristic of the detector edge regions. When using the DLA, comparable improvements are observed for gamma-ray and positron events.

### v2

@@ 1 Introduction
In the case of small-scale detectors like TAO, the time resolution of 8 ns, due to the design of the electronics system Venettacci:2024hwy , is insufficient to resolve photon time-of-flight differences. This limitation hinders the construction of precise time templates. Instead, CCA relies solely on the spatial distribution of charge to efficiently compute the charge-weighted centroid with low algorithmic complexity and high computational efficiency, making it ideal for reliable and rapid reconstruction of large event sets. DLA eliminates the need for constructing templates by directly learning nonlinear mappings from the charge and time data, achieving exceptional reconstruction accuracy. Its inherent parallel processing capabilities enable large-scale data handling, making it a powerful complement to traditional approaches. Therefore, this paper focuses on CCA and DLA and systematically investigates the optimization strategies and performance characteristics of the two algorithms, proposing the corresponding high-precision vertex reconstruction solutions for the TAO experiment and ensuring robust preparation for the upcoming data acquisition phase.
@@ 2 The TAO detector
The CD configuration of the TAO experiment is depicted in Fig. 1. The innermost component of the CD is the liquid scintillator system, which consists of a spherical acrylic vessel (AV) with an inner diameter of 1.8 $\mathrm{m}$, filled with 2.8 tonnes of GdLS. To mitigate external radioactive and incomplete energy deposition near the GdLS boundary, a fiducial volume of 0.65 $\mathrm{m}$ radius is designated as the event selection criterion.
A total of 4024 Hamamatsu S16088 Silicon Photomultiplier (SiPM) tiles, each measuring $50.7\times 50.7~\mathrm{mm}^{2}$, are installed in an array configuration on the surface of an acrylic vessel. The acrylic vessel has a thickness of $2~\mathrm{cm}$ and is supported by a copper shell structure. Each tile consists of 32 SiPMs with active areas of $12\times 6~\mathrm{mm}^{2}$, and includes two readout channels, resulting in a total of 8048 readout channels. The SiPM tile coverage in the TAO central detector surpasses 93.5%. This high coverage substantially improves photon collection efficiency, thereby ensuring an energy resolution $\sim$ 2% at 1 MeV. The data acquisition system is built upon the AD9083 ADC chip for data readout Venettacci:2024hwy , providing a 16-bit resolution and a sampling rate of $\rm 125~MS/s$. This corresponds to a time resolution of $\rm 8~ns$. In a compact detector like TAO (with a radius $\rm R<0.9~m$), this level of time resolution becomes the primary factor limiting timing measurement precision, thereby restricting the utility of timing information in the event reconstruction process.
@@ 3.2 Optimizations
First, we consider a simple case that utilizes Eq. (1) to calculate the $R$ reconstruction results for electrons directly without electronic effects. Figure 3 presents the reconstruction results for electrons with energies ranging from 1 MeV to 9 MeV, uniformly generated within the GdLS. It is observed that the reconstruction bias $B_{\text{reco}}$ becomes more pronounced near the dual-opening location, where $B_{\text{reco}}=R_{\text{reco}}-R_{\text{true}}$, with $R_{\text{reco}}$ and $R_{\text{true}}$ representing the reconstructed radius and the true radius of the energy deposition position, respectively. This is primarily due to an increase in photon leakage close to the dual-opening, resulting in a lower $R_{\text{reco}}$. To address the impact of the dual-opening effect, we propose the following dual-opening correction scheme:
@@
Fig. 6: Fitting curves of the two-dimensional distributions of $\mathit{R_{\text{true}}}$ and $\mathit{R_{\text{reco}}}$. The 137Cs curve (E = 0.662 MeV) is based on simulations along the calibration path, while the gamma ($\gamma$), positron ($e^{+}$), and electron ($e^{-}$) curves are derived from uniform simulations in the GdLS. Panels (a) and (b) show the results without and with DN correction, respectively. $R_{\rm true}$ bias” represents the difference in the $R_{\rm true}$ values between other curves and the 137Cs curve.
@@ 4.2 The VGG-T model
The Visual Geometry Group (VGG) at the University of Oxford proposed a classic CNN model known as the VGG network architecture in 2014 Simonyan:2014cmh . This model is characterized by its deep architecture, which allows for extracting hierarchical features from input data. Its advantages include improved performance on various image classification tasks, a straightforward design that facilitates transfer learning, and robustness to different image resolutions. These features make the VGG network popular for many computer vision applications. In the field of physics, the VGG network has also been widely applied. For instance, they have demonstrated promising results in pp solar neutrino identification Chen:2023xhj , dynamic aperture prediction in accelerator experiments DiCroce:2024dfy , as well as event reconstruction tasks in neutrino experiments such as PandaX Li:2022gpb , CHIPS Tingey:2022evd , and JUNO Qian:2021vnh . In this work, building upon the VGG framework, we designed a tailored VGG-T model to predict the vertex of TAO events.
@@
The predicted vertex coordinates are represented in the Cartesian space as $\mathbf{r}(x,y,z)$, with values constrained to the range [-900 mm, 900 mm]. Adopting Cartesian coordinates provides a continuous and unambiguous representation, avoiding the instabilities caused by the periodicity of angular parameters in spherical coordinates and thus simplifying the design of the loss function. To enforce this constraint, the loss function incorporates two penalty terms: $\text{Penalty}_{\text{upper}}$ for values exceeding 900 mm and $\text{Penalty}_{\text{lower}}$ for values below -900 mm. The total loss function is defined as follows:
@@ 4.3 The ResNet-T model
We utilize this technique to develop a deeper network architecture called ResNet-T, which is tailored for reconstruction tasks in the TAO experiment. As shown in Fig. 12, ResNet-T incorporates thirty-four convolutional layers beyond the original VGG architecture, resulting in a modest $1.6\%$ increase in the total number of parameters. Although the increase in the number of parameters is relatively small, the enhanced model capacity enables ResNet-T to capture more complex features and relationships in the data. Notably, this architectural modification leads to approximately a 36.5% increase in GPU memory usage, while the corresponding rise in training time is about 14.4% on an RTX-4090.
Throughout the training process, ResNet-T utilizes the same training settings as those applied to VGG-T, ensuring consistency while allowing us to explore the benefits of depth via residual connections. This approach facilitates better performance in reconstruction tasks and fosters a more stable training process, ultimately contributing to improved accuracy and efficiency in the TAO experiment. Figure 13 depicts the training dynamics of VGG-T and ResNet-T, showing the progression of the loss function and the learning rate schedule over the course of training, suggesting that both models converge smoothly without any evident signs of overfitting.
Fig. 13: Training loss and learning rate curves for the VGG-T and ResNet-T models. To improve visualization, the training and validation loss values of VGG-T are multiplied by 10 (labeled as 10×) so that their trends can be more easily distinguished from those of ResNet-T, given their originally similar magnitudes.
@@ 5 Results and discussions
This section discusses the reconstruction performance of different algorithms. Leveraging its powerful nonlinear fitting capability, DLA produces reconstruction residuals that exhibit no obvious skewness or long tails, as shown in Fig. 14 (a)–(c). However, for high-energy particles, the residuals from CCA reconstruction across the full volume exhibit slight tails, deviating from a pure Gaussian distribution, as shown in Fig. 14 (e)–(f). This phenomenon is primarily caused by increased energy leakage near the spherical shell boundary for high-energy particles, which in turn leads to more non-Gaussian deviations in the residuals. Currently, CCA cannot fully mitigate this effect. To ensure a fair and consistent comparison, we employ the mean and standard deviation of the residuals as the evaluation metrics for reconstruction bias and resolution in this work, respectively.
Fig. 14: Residual distributions within the full volume ($\rm R<900~mm$) for DLA and CCA reconstructions (in electrons). The blue histograms and red curves show the true residual distributions and their corresponding Gaussian fits, respectively. Positrons and gammas exhibit similar behavior.
@@ 5.1 Performance improvement of optimized CCA
In Section 3.2, we introduced optimizations to CCA targeting dual-opening, $R$ reconstruction nonlinearity, and electronic effects in the TAO CD to enhance reconstruction performance. Since the improvements to the reconstruction of the angles $\theta$ and $\phi$ are negligible, this section is dedicated solely to systematically evaluating and discussing the impact of these optimizations on the $R$ reconstruction.
@@ 5.2 Comparison of reconstruction performance between CCA and DLA
The reconstruction results of DLA in $R$ reconstruction are shown in Fig. 18. In terms of bias, the outcomes for all three particle types are generally controlled within 1 mm. However, positrons with initial kinetic energies below 2 MeV exhibit a larger bias exceeding 1 mm. This phenomenon arises because the gamma rays produced during the annihilation of low-energy positrons carry a higher proportion of energy, leading to a more dispersed distribution of energy deposition positions and consequently increasing the reconstruction error. Regarding reconstruction resolution, the $R$ resolution values demonstrate a decreasing trend as the deposited energy increases. Among the particles, electrons exhibit the best resolution performance, benefiting from their more concentrated energy deposition under the same deposited energy conditions, which enables more precise vertex localization and reduces statistical errors. Comparing the results of the two network architectures, ResNet-T, with its deeper network layers, shows improvements in $R$ reconstruction bias and resolution, particularly notable resolution enhancement. It is worth noting that the improvement in electron resolution is the most significant, reaching up to 13%, highlighting the substantial optimization effect of this network structure on electron reconstruction performance. In addition, we conducted both generalization and robustness tests on both models, and the test results indicate that they both meet the reconstruction accuracy requirements.
@@
Figure 19 compares electron reconstruction performance between the CCA and DLA, revealing several key advantages of the latter. The DLA yields smaller biases with more stable trends than the CCA and delivers markedly better resolution under all conditions. Although the DLA’s resolution slightly degrades within the fiducial volume (R<650R<650\,$m$), this reflects its superior handling of boundary events. These edge events, despite involving more complex electronic effects, benefit from a more concentrated charge distribution on the SiPM channels (as shown in Fig. 9(a)) and enhanced distinctiveness of image features, enabling the DLA to handle boundary events more effectively. This variation in performance suggests that DLA is particularly effective at processing spatially localized signal patterns that are characteristic of the detector edge regions. When using the DLA, comparable improvements are observed for gamma-ray and positron events.